OpenStudy (anonymous):

Differentiate 6hs+(3s^2/2){[sqrt(3)-cos(theta)]/[sin(theta)]} whereas h and s are constants.

OpenStudy (campbell_st):

so you are being asked to differentiate with respect to theta so brake the problem into its 2 parts, and rewrite it using index notation where necessary \[\frac{6hs \sqrt{3}}{2} (\sin(\theta))^{-1} - \frac{6hs}{2} \frac{\cos(\theta)}{\sin(\theta)}\] and cos/sin may be written as cot... or use the quotient rule now just apply the normal rules for differentiation.